OpenStudy (anonymous):

The weights of bats in a zoo are normally distributed with a mean of 2.2 pounds and a standard deviation of 0.3 pounds. About what percent of the bats at the zoo weigh between 1.9 pounds and 2.5 pounds? a. 34% b. 47.5% c. 81.5% d. 68% I have no clue on this one... can you please help and explain? I am COMPLETELY lost on this one this time! lol

jimthompson5910 (jim_thompson5910):

Have you heard of the empirical rule?

OpenStudy (anonymous):

nope... haha

jimthompson5910 (jim_thompson5910):

The empirical rule says that roughly 68% of the distribution is within one standard deviation of the mean

OpenStudy (anonymous):

so the answer is automatically 68% because of that??

OpenStudy (anonymous):

or no??

OpenStudy (anonymous):

i'm still confused =(

OpenStudy (anonymous):

Note that 1.9 = 2.2 -0.3 and 2.5 = 2.2+0.3 They are in the range of 1 standard deviation.  The left line indicates the 1.9 point. The middle line is the mean and is the 2.2 point the right line is the 2.5 point. in a normal distribution curve it is a rule that for the area between the left and right line is 68.2% The left and right line is found by subtracting the standard deviation from the mean and adding the standard deviation to the mean.